while riding her bike to the park and back home five times, Stacey timed herself at 60 min, 62 min, 55 min, ( the wind was helping), 58 min, and 63 min. she had set a goal of averaging 60 minutes for her rides. how many minutes will she need on her sixth ride to attain her goal?
Answer by lwsshak3(11628)

while riding her bike to the park and back home five times, Stacey timed herself at 60 min, 62 min, 55 min, ( the wind was helping), 58 min, and 63 min. she had set a goal of averaging 60 minutes for her rides. how many minutes will she need on her sixth ride to attain her goal?
***
let x=travel time of her sixth ride
(60+62+55+58+63+x)/6=60
298+x=6*60=360
x=360-298=62
how many minutes will she need on her sixth ride? 62